### v2

# Benchmarking General Purpose In-Context Learning

Abstract
In-context learning (ICL) capabilities are becoming increasingly appealing for building general intelligence due to their sample efficiency and independence from artificial optimization skills. To enhance generalization, biological neural systems primarily inherit learning capabilities and subsequently refine their memory, acquiring diverse skills and knowledge through extensive lifelong experiences. This process gives rise to the concept of general-purpose in-context learning (GPICL). Compared to standard ICL, GPICL addresses a broader range of tasks, extends learning horizons, and starts at a lower zero-shot baseline. We introduce two lightweight but insightful benchmarks specifically crafted to train and evaluate GPICL functionalities. Each benchmark includes a vast number of tasks characterized by significant task variance and minimal transferable knowledge among tasks, facilitating lifelong in-context learning through continuous generation and interaction. These features pose significant challenges for models that rely on context or interactions to improve their proficiency, including language models, decision models, and world models. Our experiments reveal that parameter scale alone may not be crucial for ICL or GPICL, suggesting alternative approaches such as increasing the scale of contexts and memory states.
00footnotetext: Works done at Baidu Inc. and University of Science and Technology of China.00footnotetext: Corresponds to: Fan Wang (fanwang.px@gmail.com) and Yang Cao (forrest@ustc.edu.cn)00footnotetext: source code available at https://github.com/FutureAGI/L3C
1 Introduction
The success of large-scale generative language models can be primarily attributed to two key factors: 1. Zero-shot generalization derived from the extensive accumulation and storage of knowledge within the model parameters during the pretraining and fine-tuning stages (Kojima et al., 2022); 2. In-context learning(ICL) to distill knowledge from the context during the inference stage (Brown et al., 2020). The ability of zero-shot generalization remains unchanged post-training, which parallel the concept of “innate abilities” (Koulakov et al., 2021) of biology. The ICL capability allows the agent to handle various unseen tasks without modifying the model parameters, only resulting in changes to the hidden states and memories. Currently, the ICL capabilities of large language models are confined to tasks that solely involve interpreting natural language tokens and following instructions (i.e., mimicking demonstrations). However, they generally lack the ability to continually interact with real-world environments and engage in reinforcement learning from contexts. Additionally, extremely complex tasks that require substantial amounts of context still pose significant challenges.
Methodology Zero-shot Capability ICL
Potential
ICL
Horizon
Multi-task Learning Medium Low Short
Meta-Learning Low Medium Short
ICL High Medium Medium
GPICL Low High Long
Figure 1: Conceptual diagram of in-context learning (ICL) and general purpose in-context learning (GPICL)
The community has long recognized a significant gap between artificial intelligence and biological intelligence, particularly regarding innate abilities and the functionality to continuously learn throughout their lifespans (Parisi et al., 2019). For example, the initial limited capabilities of mammal infants stand in stark contrast to their skills in lifelong learning and adaptation. This disparity is further highlighted by the phenomenon known as the genomics bottleneck. Understanding this could guide the design of machine intelligence systems (Zador, 2019; Kirsch and Schmidhuber, 2021; Wang et al., 2022). A promising research area is general purpose in-context learning (GPICL) (Kirsch et al., 2022; Wang et al., 2022). GPICL involves meta-training across a wide range of task classes and learning to interpret context within a variety of highly diverse tasks. However, the characteristic of diversity in this context can be challenging to quantify. We aim to enrich the concept of GPICL by incorporating additional features, as shown in Figure 2. Consider an agent that ideally utilizes all available data, including meta-training data and contextual information; The GPICL system should corresponds to low zero-shot capability, high ICL potential, and an extended ICL horizon.
(a) LLM mainstream learning pipeline
(b) GPICL learning pipeline
Figure 2: Learning pipelines based on GPICL versus the current LLM learning pipeline (Ouyang et al., 2022).
Why GPICL? The objective of extending the in-context learning (ICL) horizon and potential is widely acknowledged. However, it raises the question of why we should reduce zero-shot capability, which appears to contradict the prevailing trend in large language models. First, inspired by the genomics bottleneck and building on previous work (Kirsch and Schmidhuber, 2021), one potential strategy to enhance generalizability of the meta-model is to maintain compact meta-parameters that are focused primarily on the functionality of learning itself, rather than to keep overwhelming parameters designed to capture all task-specific knowledge and skills. Second, the use of abundant synthetic data in pre-training and meta-training, which tends to be non-factual and biased, presents challenges to the usage of the model. In this context, GPICL could offer an ideal solution: since the meta-training stage can result in the loss of much task-specific knowledge and skills but preserves the core learning functionality, it helps prevent the meta-model from being overly influenced by non-factual information and bias in those tasks. Third, in practice, a model’s performance typically depends on both in-context learning (ICL) and zero-shot generalization capabilities, making it challenging to investigate and improve ICL independently. Isolating the ICL capability alone would be beneficial for inspiring better model architectures and learning algorithms.
Building on those points, there can be an prospective approach for advancing the development of general intelligence in contemporary settings: Initially, a substantial volume of synthesized data, potentially containing a wealth of nonfactual and biased information, is used exclusively to cultivate the GPICL functionality. Subsequently, a considerable amount of high-quality, unbiased, well-aligned data is employed to enrich the contexts and memory, enabling the agent to achieve capabilities approaching those of humans. A visual illustration of this proposed pipeline with existing methodologies is depicted in Figure 2.
Guidelines for designing GPICL benchmarks. We note that a major impediment to research in this area is the lack of benchmarks that adhere to the GPICL standard. We suggest that those benchmarks should satisfy the following standard:
  • Vast number of variant tasks with minimal knowledge transfer: To minimize the effectiveness of zero-shot generalization, it is advantageous to engage with a variety of tasks where task-specific knowledge cannot be transferred among different tasks. This guarantees that the model’s performance is primarily driven by its in-context learning capabilities rather than its ability of zero-shot generalization. It is essential to require numerous tasks to prevent the meta-model from hacking in-context learning (ICL) by merely memorizing all possible tasks and performing task identification, a limitation that most meta-learning benchmarks fall into when given limited tasks.
  • Interactive and generative tasks: GPICL should include task exploration and completion processes that are both interactive and generative, such that in-context reinforcement learning capabilities can be effectively meta-trained and evaluated. Consequently, benchmarks that solely focus on encoding a single sample characterized by extended contexts and features and conducting single-step predictions or classifications do not meet the criteria.
  • Lifelong in-context learning: GPICL should handle very long-term dependencies that can scale up to millions or even billions of steps, effectively supporting the learning process over extensive periods (ICL horizon $T$ should be “from cradle to grave”). It is necessary to clarify that $T$ should denote the “minimum steps” possible given “ideal ICL functionality”. While many previous gradient-based meta-learning methods also encode adaptation over millions of samples (e.g., many gradient-based meta-learning methods), most of these are to blame on the inefficient learning functionality. Consequently, they should not be considered as representing a long ICL horizon.
Contribution of this paper: We propose two critical benchmarks that adhere to the GPICL standard and are capable of generating unlimited synthetic data: 1. Meta-Language: We create a randomized yet consistent language pattern to meta-train models, demonstrating that a model, without prior exposure to natural language corpora, can acquire the ability to learn a new language from scratch through ICL; 2. Maze World: We generate an unlimited number of tasks within randomized mazes, where the model is meta-trained solely through context to learn exploration and navigation. For both benchmarks, we provide limited experiment results of baseline models based on Transformer (Vaswani et al., 2017) backbone. The primary goal of this paper is neither to introduce a new model or learning methodology, nor to provide an exhaustive survey of the existing ICL systems; rather, we aim to demonstrate through these baselines the potential for these benchmarks to benefit the research of ICL and GPICL. One crucial conclusion from these limited experiments may be that parameter scale alone is not crucial for ICL or GPICL. This finding aligns with previous research indicating that the size of contexts and memory states can play a more significant role (Kirsch et al., 2022; Wang et al., 2022). This insight may highlight alternative approaches beyond merely scaling up models within the current AI community.
2 Learning to learn an randomized language: Meta-Language
2.1 Problem Setting
Most large language models (LLMs) are trained on specific, existing languages. Rather than synthesizing existing languages, we primarily consider learning a new language as a task to focus more on the ICL functionality. Although there are hundreds of languages in the world, treating each language as a sample for GPICL provides only a few hundred samples, which is insufficient. We introduce the "meta language" benchmark. This benchmark is designed to generate a vast number of new "languages" and assess the models’ ability to learn an arbitrary new "language." This pseudo-language is simulated using n-gram settings, generated by a randomized neural network $f$. The probability of the next token is determined by its preceding n tokens and a set of random parameters $\theta$.
$p(x_{t}|x_{t-1},x_{t-2},...)=f(x_{t-1},...,x_{t-n};\theta),$ (1)
Algorithm 1 Generating a length=T sequence of meta langauge
1:Randomly Sample $\theta\sim\mathcal{N}(0,\sigma^{2})$, ($\theta$ is the parameter of the specified n-gram model $f$)
2:for t in [1, T] do
3:     $z_{1},...,z_{N}=f(x_{t-1},...,x_{t-n};\theta)$
4:     For $i\in[1,N]$, $z_{i}=\lambda[z_{i}-\mathbb{E}(z)]/\sqrt{\mathbb{V}(z)}$
5:     $p(x_{t}=x_{i})={exp(z_{i})}/{\sum_{j}exp(z_{j})}$
6:     Randomly sample $x_{t}\sim p(x_{t})$
The generation process is as simple as Algorithm 1. To maintain the diversity of the generated meta language, we also normalize the mean and variance of the logits input to Softmax to keep the distribution being neither too stiff nor too flat. We set the hyper-parameter $\lambda$ in Algorithm 1 such that the perplexity of the ground truth generator $\mathbb{E}[-logp(x_{t}))]$ is kept between $0.5$ and $1.0$. The sequences generated by the randomized generator are meaningless and chaotic. However, as long as the generator’s parameters remain static, a model can progressively learn and capture the underlying rules of the randomized generator. The complexity of the generated sequences can be bounded by $N^{n}$, where $N$ is the size of the vocabulary. Therefore, with a sufficiently large $N$ and $n$—parameters that can be adjusted as needed—the complexity of the pseudo-language can be made arbitrarily high to test super-long-term dependencies.
2.2 Baseline Solution
As the training data, we generated 500K sequences with a pre-defined vocabulary size of $N=32$, and the complexity of the pseudo-language was uniformly sampled across $n\in{3,4,5,6}$. Each sequence comprises $4,096$ tokens, totaling 2B tokens. We apply the auto-regressive transformer (Vaswani et al., 2017) with rotary embedding (Su et al., 2024). The hyper-parameter settings include: a tiny-sized transformer (303k parameters), a small-sized transformer (9.5m parameters), and a standard-sized(151m parameters) model. The evaluation set consists of independently sampled sequences, totaling 4K samples (16M tokens) for each $n\in\{2,3,4,5,6,7,8\}$, sampled uniformly. We assessed the model’s performance by plotting the averaged perplexity ($-\log p(x_{t})$) for each position across different $n$ values. To assess how the trained model generalizes, we also included the PG-19 (Rae et al., 2019) test set as a natural corpus. In this set, words are broken down into 26 alphabetic characters and randomly mapped onto 32 pre-defined tokens, the left slot are given to punctuation marks. The entire test texts are concatenated into a long sequence and broken down into sentences of 4,096 characters each. It is important to note that the model has never been trained on any natural language corpora. Specific details are provided in the appendices.
2.3 Evaluation Results
2.3.1 Towards Meta-Language Model
Figure 3: Perplexity per position on evaluation data sampled by different language complexity.
We demonstrate that this pseudo-language trained model can be regarded as a meta-language model designed to adapt to any conceivable language through GPICL. See from basic evaluation results of 151M model on the test sets in Figure 3, several observations are noteworthy: First, perplexity improves with increasing context length across all values of $n$, indicating that the model is effectively leveraging in-context learning. Notably, for $n=7$ and $n=8$—complexities on which the model was not explicitly trained—we observed a consistent semi-logarithmic improvement in perplexity with increasing context length. Second, a higher language complexity, indicated by larger values of $n$, correlates with a more gradual improvement in perplexity relative to context length. This phenomenon confirms that increased language complexity introduces longer-term dependencies within contexts, establishing meta language as an ideal benchmark for evaluating long-dependency modeling capabilities. Most importantly, the model demonstrates a significant performance gain through in-context learning (ICL) in realistic natural language scenarios, despite never being exposed to a single real-world natural word, illustrating that the ICL capability acquired through meta-language is not restricted to a specific language type.
We test the ability of generation of the meta-language model on two rudimentary real-world language tasks: English vocabulary learning and basic mathematical operations (addition and subtraction), which are typical for primary school education. For the English vocabulary task, the vocabulary consists of alphabet characters and common punctuation marks; for the math task, it includes digits and mathematical operators. The input contexts for testing included 40 English words and simple arithmetic operations below $5$. To discourage direct replication, we intentionally introduced errors into $15\%$ of the words or equations. We analyzed the outputs of the MetaLM across two different context lengths, with the shorter contexts comprising only the first half of the longer contexts. The results, depicted in Fig. 4, show that MetaLM is capable of learning to memorize and correct both English vocabulary and mathematical equations, validating that its ICL capability is independent of vocabulary settings. Furthermore, the longer the context provided, the fewer errors appeared in the outputs, indicating a clear trend of improvement.
(a) English vocabulary learning task.
(b) Mathematical operation task.
Figure 4: Input contexts were designed such that $15\%$ contained errors. All mistakes are highlighted with red squares. The vocabulary for both tasks was limited to 32 pre-defined tokens, and there was no pre-training or further tuning applied to either task. Notice that this setup tests the MetaLM’s ability to adapt and correct based on in-context learning alone, without relying on prior knowledge or adjustments. The model under evaluation has not been trained on any natural langauge corpora. Additionally, the vocabulary is randomly assigned to the predefined 32 tokens prior to inference.
2.3.2 Scaling Laws for GPICL
We have conducted further research into how the scale of parameters affects GPICL capability. As illustrated in Fig. 5, our findings reveal notable trends: 1. Language complexity ($n$) influences the relative performance of different-sized models and affects the dependence on context length. In simpler tasks ($n<4$), performance rapidly converges to lower bounds asymptotically ($t\sim T=4k$. Here, with a slight abuse of notation, $T$ does not necessarily represent the ICL horizon, but rather denotes the maximum context length used in our experiments.), and the performances of various models become indistinguishable, despite starting differently. 2. It should be noted that for tasks as straightforward as $n=2$, where a 300K model achieves comparable asymptotic performance to a 150M model, we observe that larger models cannot attain higher performance without sufficient context. It is reasonable to infer that increasing model complexity further will not be beneficial. This observation aligns with a key characteristic of GPICL, where learning from context takes precedence over zero-shot generalization. 3. For greater complexity ($n>6$), the ICL horizon clearly extends far beyond $T=4k$, and the improvement in perplexity exhibits polynomial scaling law (depicted as a linear relationship on a semi-logarithmic axis) with increasing context length. However, it remains difficult to assert from current experiments whether, given sufficiently long contexts, models of varying scales will converge to similar asymptotic performances. Nevertheless, we can confidently conclude that for $n<8$, a model with 10 million parameters has been adequate. 4. For the PG-19 test set, we were surprised to find that the tiny-sized model achieves comparable asymptotic performance to the other two models. We must consider that the complexity of natural language far exceeds that of any conceivable $n$-gram language. Furthermore, the improvement does not exhibit a linear trend on a semi-logarithmic scale. It is reasonable to hypothesize that the GPICL is only capturing certain short-term and high-frequency patterns in natural language (such as frequently appeared words).
Our experiments suggest a shift in the GPICL paradigm from “the larger the parameters, the better” to “the longer the context (and the larger the memory), the better”, provided the parameters are sufficient. In contrast, current large language models (LLMs) typically require at least billions of parameters, the majority of which we suspect are employed for zero-shot generalization, while the parameters dedicated to Incremental Context Learning (ICL) are relatively sparse. This perspective is consistent with the concept of the genomics bottleneck (Wang et al., 2022), which advocates for models that, while adequately parameterized, prioritize larger memory capacities to optimize GPICL. However, it is important to note that while the transformer architecture offers a “full memory” capability for sequence modeling, it also incurs a computational cost of $\Theta(T^{2})$. This underscores the necessity for developing more efficient transformers, which would allow for the extension of context length to longer spans. We believe research in this area (Tay et al., 2022) could benefit from this benchmark.
Figure 5: Evaluating language models of various sizes on Meta-Language benchmark with differing levels of complexity ($n$).
3 Learning to navigate an randomized world: Maze World
3.1 Problem Setting
The Maze World simulation tasks an agent with navigating a completely unknown environment, requiring it to explore, memorize its surroundings, perform simultaneous localization and mapping (SLAM), identify navigation targets, and plan its route. This setting could potentially reflect a wide range of real-world tasks, such as household robots adapting to unfamiliar indoor environments. As the agent gains more experience within a specific environment, it should progressively gather more information and navigate to specified targets more efficiently. The proposed benchmarks has the following features:
  • Each task corresponds to a randomly generated $n\times n$ maze, with layouts and textures assigned randomly.
  • A number of Potential Navigation Targets (PNTs) are placed at random locations and remain fixed throughout the task. Each PNT is marked by a semi-transparent beam of light, as shown in Figure 6(a).
  • For each task, a randomized sequence of commands is generated, each directing the agent to walk to a specified target that is identifiable by color.
The agent receives a positive reward for reaching the specified target, while a minor cost is incurred for each step taken. Available actions include moving forward, moving backward, stopping, turning left, and turning right. Beyond these task settings, the benchmarks provide a wide range of hyper-parameters to generate diverse tasks. Examples include the height and angle of the perspective, ceiling height, and the length of the cell unit. In addition to command-based navigation, we also offer a “Survival” mode, details are provided in the appendices.
To optimize the initiation of model training, we have implemented rule-based smart agents, which are allowed to access the global map, but with limitations. They record everything within their line of sight into memories, while areas out of sight remain hidden. This kind of agent efficiently navigates to its goals by utilizing an exploration-then-exploitation strategy. Although this rule-based smart agent may not consistently reach the global optimum, primarily due to shortcomings in its exploration strategy, it nonetheless provides a solid baseline for imitation learning.
To model varied agent behaviors, we differentiate between Long-Term Memory (LTM) and Short-Term Memory (STM) in the rule-based agents. The STM retains only the most recent three steps of observations, while the LTM retains information until the episode concludes. We introduce a hyper-parameter, $p(\text{STM}\rightarrow\text{LTM})$, to govern the probability of transferring information from STM to LTM. At $p(\text{STM}\rightarrow\text{LTM})=0\%$, the agent operates without LTM, whereas $p(\text{STM}\rightarrow\text{LTM})=100\%$ allows for a fully functional LTM.
(a) A demonstration of observations and a global map in MazeWorld
(b) A sketch of an agent’s LTM
Figure 6: A demonstration of the maze world tasks alongside a rule-based navigation system. The rule-based agent is provided with partial access to the environment’s ground truth (b) under specific conditions. Only the areas within the agent’s line of sight are revealed, while the unsighted regions remain obscured. The agent retains the visual information from the nearest three frames in its short-term memory, which is then transferred to long-term memory based on a specified probability, denoted as $p(\textit{STM}\rightarrow\textit{LTM})$ The agent employs a strategy that balances exploration with exploitation.
3.2 Baseline Solution
We offer a baseline solution for the Maze World tasks, which is largely inspired by model-based imitation learning (Hu et al., 2022), causal decision models (Chen et al., 2021), and meta-reinforcement learning (Duan et al., 2016). We use Variational Auto Encoder (VAE) to process each image to latent representation, and use a causal transformer (Chen et al., 2021) to encode the historical observation and action. The model ouputs the distribution of actions, and also the expected next frame. Specifically, given the reference trajectory $(i_{1},a_{1},i_{2},a_{2},...,i_{T},a_{T},i_{T+1})$, where $i_{t}$ and $a_{t}$ represent the observation and action respectively, our model is represented by:
$z_{t}$ $=\textbf{ENC}_{\text{vae}}(i_{t}),$ (2)
$\hat{i}_{t}$ $=\textbf{DEC}_{\text{vae}}(z_{t}),$ (3)
$\pi(a_{0}),\hat{z}_{1},\pi(a_{1}),\hat{z}_{2},...$ $=\textbf{CAUSAL}(z_{0},a_{0},z_{1},a_{1},...),$ (4)
where $z_{t}$ is the latent embedding of the observed image $i_{t}$. Predicting the distribution of actions ($\pi$) and the next frame ($\hat{z}_{t}$) are typically referred as Policy Model and World Model respectively.
We train the model using imitation learning, although reinforcement learning is desirable for future enhancements. The losses are composed of three parts: 1. the reconstruction error ($\mathcal{L}_{vae}$); 2. the cross entropy between the reference action and the predicted distribution ($\mathcal{L}_{pm}$); 3. the next-frame prediction error ($\mathcal{L}_{wm}$). The final loss would be:
$\mathcal{L}_{vae}$ $=\sum_{t}||\hat{i}_{t}-i_{t}||^{2}+\lambda\text{KL}(p(z_{t})||\mathcal{N}(0,1))$ (5)
$\mathcal{L}_{pm}$ $=\sum_{t}-log\pi(a_{t})$ (6)
$\mathcal{L}_{wm}$ $=\sum_{t}||\textbf{DEC}_{\text{vae}}(\hat{z}_{t})-i_{t}||^{2}$ (7)
$\mathcal{L}$ $=\alpha_{1}\mathcal{L}_{vae}+\alpha_{2}\mathcal{L}_{pm}+\alpha_{3}\mathcal{L}_{wm}$ (8)
The choice of causal transformer includes a small-sized transformer (26M parameters) and a standard-sized transformer (330M parameters). A overview of the model architecture is in Fig. 7. To train our baseline models, we collect training data include 12.8K episodes of demonstration from rule agents for imitation learning. To avoid the growing “compounding errors” (Ross and Bagnell, 2010) associated with direct behavior cloning, we drew inspiration from data augmentation methods (Ross et al., 2011). We collected the trajectory of states by running a very noisy behavior policy, but used a rule-based agent to label its action at each step (which is not actually executed; see appendices for details). Each episode has $2,048$ steps, which gives $26$M frames in total.
Figure 7: An overview of the model evaluated on the maze world tasks. Initially, a Variational Auto-Encoder (VAE) is trained to encode observed images into hidden vectors and vice versa, with its parameters subsequently frozen. The sequence of observations and actions is then utilized to train a causal model, specifically a transformer that employs only backward attention, using imitation learning and self-supervised learning techniques.
3.3 Baseline Performances
(a) $\mathcal{L}_{wm}(t)$
(b) $\mathcal{L}_{pm}(t)$
Figure 8: Static assessment of the ability to mimic rule-based agents ($\mathcal{L}_{pm}$) and predict the next frame ($\mathcal{L}_{wm}$) across each time step ($t$). Context lengths of 512 and 2048 are evaluated respectively for comparison.
Static Evaluation. We first conduct static evaluation using off-policy data from an independently sampled validation set comprising $720$ sequences (totaling $1.5$M frames). To analyze the impact of context length on performance, we assessed the losses of the world model and policy model for each individual time step ($\mathcal{L}_{wm}(t)$ and $\mathcal{L}_{pm}(t)$). In Figure 8, We compare the performance of the small-sized and standard-sized causal models, with several noteworthy observations: First, the future prediction task shows no clear dependency on long contexts (the learning horizon is below 100 steps), whereas the performance of action prediction continually improves with longer contexts. This likely suggests that world modeling in the proposed tasks is more related to few-shot ICL capabilities than to GPICL. This is unsurprising, since single-step world modeling is more connected with understanding the localized environment, whereas navigation requires more global reasoning. Second, the standard-sized model does not demonstrate a clear advantage over the small-sized model. This phenomenon aligns with that observed in meta-language tasks, further validating the idea that scaling up parameters is not always necessary compared to scaling up contexts in GPICL.
(a) Maze $15\times 15$
(b) Maze $25\times 25$
(c) Maze $35\times 35$
Figure 9: Interactive evaluation results of the rule-based agents, randomized policy, and the causal model, depicted by plotting accumulated rewards against simulated time steps. Each plot summarizes the results from identical 32 tasks. The step minor cost is set to 0.
Interactive Evaluation. It is more beneficial to investigate the online performance of the policy model and world model in causal modeling by directly interacting with the environment.
We first directly evaluate the model-free policy by interacting with a set of evaluation tasks. The evaluation set includes 32 tasks across mazes of sizes $15\times 15$, $25\times 25$, $35\times 35$. A softmax policy, based on the predicted distribution of actions, is implemented. We plot the mean and variance of the average accumulated rewards over time (t). For comparative analysis, we also include the performance of a random policy, where actions are randomly sampled regardless of the observations, and agents with different short-term memory to long-term memory transition probabilities $p(\text{STM}\rightarrow\text{LTM})$ of $5\%$, $25\%$, $100\%$ (Figure 9). The results indicate that our model significantly outperforms random exploration and rule-based agents with low long-term memory capabilities ($p(\text{STM}\rightarrow\text{LTM})=5\%$). However, there is a noticeable gap compared to those with more effective long-term memories ($p(\text{STM}\rightarrow\text{LTM})=25\%,100\%$). Moreover, the rate of improvement is not significant compared to those with effective long-term memories, where we would expect to see an accelerating accumulation of rewards. We believe that there is ample room for improvement by employing more sophisticated models and better learning algorithms, and extending the context length.
Figure 10: Randomly selected clips depicting the future 8 steps using the world model of causal modeling are presented. For each clip, the upper section shows the ground truth observations from the current step to 8 steps ahead (from left to right), while the bottom section shows the imagined future up to 8 steps ahead starting from the current observation. The actions taken are also labeled with text in the image. The future imagination is generally accurate for the first frame but deteriorates over time due to error propagation, illustrating the challenge of causal world modeling.
We then provide a qualitative evaluation of the performance of world models in Figure 10. Unlike $\mathcal{L}_{wm}$, which can only represent the precision of predicting one-step transitions, using the world model to predict the distant future is far more challenging. We observe that although our world model can predict the nearest future, its accuracy quickly deteriorates over time. This can be attributed to the well-known problem of horizon-square compounding errors. Typical improvements include adding noise to the inputs (Hu et al., 2022) and multi-step prediction (Asadi et al., 2019). However, compared to many previous problems, Maze World poses a higher challenge to world modeling in two key aspects: First, the variance in observations is more severe (the observation of two consecutive frames can be very different). Second, the variance of the tasks requires the world model to perform in-context learning to improve, as each maze is completely different and not transferrable.
4 Related Work
4.1 Meta-learning and GPICL
General purpose in-context learning (GPICL) is closely related to meta-learning , particularly model-based meta-learning (Duan et al., 2016; Santoro et al., 2016). The latter generally encompasses a wide range of methodologies, including those that rely on gradient-based adaptation (Finn et al., 2017), while in-context learning focuses on gradient-free “instruction following” (Ouyang et al., 2022). In recent years, the distinction between meta-learning and pre-training-ICL has become increasingly blurred (Reed et al., 2022; Dai et al., 2023; Li et al., 2023). However, GPICL fundamentally differs from traditional model-based meta-learning and ICL in large-scale language models in the key aspects including minimal zero-shot generalization, long learning horizon, and abundancy of learning algorithm (where supervised learning, reinforcement learning and others are included).
4.2 Benchmarks for Meta-Learning and ICL
A variety of benchmarks have been utilized to evaluate meta-learning algorithms, encompassing domains such as image classification (LeCun et al., 1998), video games, and robot manipulation (Brockman et al., 2016). Many of these benchmarks were not initially designed for meta-learning but have been adapted to accommodate varying tasks through parametric adjustments in hyper-parameters, such as target positions and robot configurations. Meta-World (Yu et al., 2020) introduces a series of robot arm manipulation tasks that incorporate both parametric and non-parametric variations. Mazes are also frequently used to benchmark meta-reinforcement learning (Duan et al., 2016; Mishra et al., 2018), but the scale and diversity of those tasks are not adequate enough to benchmark GPICL.
A key distinction between our proposed benchmarks and previous ones in meta-learning is that the latter often veer towards "task identification" by selectively omitting critical observations. For example, scenarios involve a robot moving in a direction that remains hidden and can only be inferred from rewards (Finn et al., 2017), or a robot arm reaching for a target that is neither visible nor directly observed (Yu et al., 2020). In practical settings, such information like targets and directions would typically be explicitly indicated through instructions or visual cues, making the tasks ostensibly complex yet fundamentally simplistic. Moreover, these tasks generally do not necessitate long-term interaction, with most being solvable through relatively brief explorations. Our benchmarks, in contrast, align more closely with real-world challenges. For instance, the Maze World “navigation” tasks can mirror the adaptive learning process of a household robot adapting to a new indoor scenario, where it needs to simultaneously solve localization, mapping, and planning. Another group of Maze World “survival” tasks simulate an agent exploring a new environment to learn about helpful and harmful elements.
Another set of benchmarks has emerged alongside large language models (LLMs), focusing on evaluating the capability of modeling long sequences. The Long Range Arena (Tay et al., 2020) includes six tasks, primarily adapting image classification tasks to pixel-wise classification. Other benchmarks involve natural corpora, such as literary texts (Rae et al., 2019) and long document classification (Chalkidis et al., 2019). However, a significant issue with using these benchmarks to study GPICL is that it is often difficult to distinguish between GPICL capabilities and zero-shot generalization. Indeed, previous research on efficient transformers (Tay et al., 2022) has shown that improvements in zero-shot generalization (such as through simply increasing model scale) tend to outperform increases in context length (especially beyond 4,000 tokens) in these datasets. This suggests that these benchmarks may not be ideal for thoroughly testing GPICL.
5 Conclusions
In this paper, we enrich and redefine the topic of General Purpose In-Context Learning (GPICL) and underscore its significance to the AI community. We introduce two benchmarks specifically designed for GPICL. Our preliminary investigations of these benchmarks have yielded some inspiring results, demonstrating that a completely randomized meta-language can be used to train a language model capable of understanding realistic natural language to a certain extent.
It is important to note that although we are seeing the potential of application of those datasets in areas including language modeling and indoor navigation, the synthetic data generated by those benchmarks are not yet validated to be ready for a direct help to application. Currently, large language models (LLMs) typically feature vocabulary sizes $N$ exceeding tens of thousands, highlighting that the true complexity of natural language far surpasses $8$-gram. Moreover, the visual signals in maze worlds remain regularized and relatively monotonic, still significantly less complex than those found in the real world.
This work can be expanded along two directions: 1. The development of more realistic GPICL benchmarks, for instance, by adding greater complexity to the meta-language; 2. Utilizing the benchmarks to uncover more sophisticated models (including memory-augmented models) and optimization techniques (such as reinforcement learning) to further extend the GPICL horizon and enhance performance gains.